The volumes of soda in a quart soda bottles are normally distributed with a mean of 32.3 oz and a standard deviation of 1.2 oz. What is the probability that the volume of soda in a randomly selected bottle will be less than 33 oz?
Answer by stanbon(75887)

The volumes of soda in a quart soda bottles are normally distributed with a mean of 32.3 oz and a standard deviation of 1.2 oz. What is the probability that the volume of soda in a randomly selected bottle will be less than 33 oz?
------
z(33) = (33-32.3)/1.2 = 0.58333..
------------
P(x < 33) = P(z < 0.58333..) = normalcdf(-100,0.58333) = 0.7202
---
Cheers,
Stan H.
--------